solve by factoring and using principle of zero product.
5x^2-66=4x^2+5
Am I starting it off correctly
5x^2-66=0 and 4x^2+5=0 then solve for x^2?




Answer by Alan3354(69443)

5x^2-66=4x^2+5
Am I starting it off correctly
5x^2-66=0 and 4x^2+5=0 then solve for x^2?
------------------
You can't say they're equal to zero, they're not.
----------------
5x^2-66=4x^2+5
x^2 = 71
x = +/- sqrt(71)
-----------------
I don't see anything to factor.